OpenStudy (anonymous):

minimize x^2+y^2 if 2x+y=10

OpenStudy (gebooors):

First solve y = 10-2x, then put it into equation, you get X^2 + (10-2x)^2 = x^2 + (100 -40x + 4x^2)

OpenStudy (anonymous):

looks like i solved the question here @kyle2393 ----------------->>> http://openstudy.com/study#/updates/5336ec65e4b0b47fca85e0ed

OpenStudy (gebooors):

I also saw same question in mathematics, Kyle should give you medal

OpenStudy (anonymous):

well... i donot want a medal @Gebooors .... the point is it that s/he posted the same question at many places and thats what i wanted to say Kayle

OpenStudy (gebooors):

We all do mistakes. Don't worry Kyle:)